OpenStudy (rsadhvika):

Is there a relationship between decimal expansion of a number and repeated partial fractions ?

OpenStudy (rsadhvika):

\[\dfrac{x^{n-1}}{(x-a)^n} = \dfrac{A}{x-a} + \dfrac{B}{(x-a)^2} + \dfrac{C}{(x-a)^3}+\cdots\]

OpenStudy (ikram002p):

decimal expansion ?

OpenStudy (rsadhvika):

yes

OpenStudy (ikram002p):

like what do u mean by it xD

OpenStudy (rsadhvika):

\[0.\overline{123} = \dfrac{1}{10} + \dfrac{2}{10^2} + \dfrac{3}{10^3} + \cdots \]